OpenStudy (anonymous):

find x and y intercepts y=-2/3x + 1/3 how do i do this?

OpenStudy (anonymous):

X-intercept: when y=0 Y-intercept: when x=0

OpenStudy (anonymous):

i got 1/18 for x but i think that wrong!?

OpenStudy (anonymous):

Put 0 where y is and solve for x. 0=-2/3x+1/3 -1/3=-2/3x 1/2=x

OpenStudy (anonymous):

you didvide both by -2/3?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

till dont get it when i divide-1/3 by the -2/3 i still get 1/18???

OpenStudy (anonymous):

When dividing by a fraction, you flip the fraction and instead multiply. -1/3*-3/2

OpenStudy (anonymous):

oh ok thanks so much :)

OpenStudy (anonymous):

So x intercept is (1/2,0) what did you get for y intercept?

OpenStudy (anonymous):

i got (0,1/3) ?